# Short-term CO2 emissions forecasting: insight from the Italian electricity market

### v1

@@ Appendix B Tables
SARIMAX(3,1,0)(1,0,0)24 FARX(1) GAM(3,1) TBATS KNN(10) COMB2.BG COMB2.SEL COMB4.SEL
SARIMAX(3,1,0)(1,0,0)24 $<0.05$ $<0.05$ 1.0000 0.8817 $<0.05$ $<0.05$ $<0.05$
FARX(1) 1.$0$ $<0.05$ 1.0000 1.0000 $<0.05$ $<0.05$ $<0.05$
GAM(3,1) 1.0000 1.$0$ 1.0000 1.0000 0.8872 $<0.05$ $<0.05$
TBATS $<0.05$ $<0.05$ $<0.05$ $<0.05$ $<0.05$ $<0.05$ $<0.05$
KNN(10) 0.1183 $<0.05$ $<0.05$ 1.$0$ $<0.05$ $<0.05$ $<0.05$
COMB2.BG 1.0000 1.0000 0.1128 1.0000 1.$0$ $<0.05$ $<0.05$
COMB2.SEL 1.0000 1.0000 1.0000 1.0000 1.0000 1.$0$ 1.0000
COMB4.SEL 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 $<0.05$
@@
SARIMA(1,1,0)(1,0,0)24 FARX(3) GAM(3,1) TBATS KNN(10) COMB2.BG COMB2.SEL COMB4.SEL
SARIMA(1,1,0)(1,0,0)24 $<0.05$ $<0.05$ 1.0000 $<0.05$ $<0.05$ $<0.05$ $<0.05$
FARX(3) 1.$0$ $<0.05$ 1.0000 1.0000 $<0.05$ $<0.05$ $<0.05$
GAM(3,1) 1.0000 1.$0$ 1.0000 1.0000 $<0.05$ $<0.05$ $<0.05$
TBATS $<0.05$ $<0.05$ $<0.05$ $<0.05$ $<0.05$ $<0.05$ $<0.05$
KNN(10) 1.0000 $<0.05$ $<0.05$ 1.$0$ $<0.05$ $<0.05$ $<0.05$
COMB2.BG 1.0000 1.0000 1.0000 1.0000 1.$0$ $<0.05$ 0.4581
COMB2.SEL 1.0000 1.0000 1.0000 1.0000 1.0000 1.$0$ 1.0000
COMB4.SEL 1.0000 1.0000 1.0000 1.0000 1.0000 0.5419 $<0.05$
@@
SSARIMAX(3,1,0)(1,0,0)24 FAR(2) GAM(3,1) TBATS KNN(15) COMB2.BG COMB2.SEL COMB4.SEL
SARIMAX(3,1,0)(1,0,0)24 - $<0.05$ $<0.05$ $<0.05$ $<0.05$ $<0.05$ $<0.05$ $<0.05$
FAR(2) 1.$0$ $<0.05$ 1.0000 1.0000 $<0.05$ $<0.05$ $<0.05$
GAM(3,1) 1.0000 1.$0$ 1.0000 1.0000 $<0.05$ $<0.05$ $<0.05$
TBATS 1.0000 $<0.05$ $<0.05$ $<0.05$ $<0.05$ $<0.05$ $<0.05$
KNN(15) 1.0000 $<0.05$ $<0.05$ 1.$0$ $<0.05$ $<0.05$ $<0.05$
COMB2.BG 1.0000 1.0000 1.0000 1.0000 1.$0$ $<0.05$ 0.9631
COMB2.SEL 1.0000 1.0000 1.0000 1.0000 1.0000 1.$0$ 1.0000
COMB4.SEL 1.0000 1.0000 0.9999 1.0000 1.0000 $<0.05$ $<0.05$
@@
SARIMAX(2,1,0)(1,0,0)24 FARX(3) GAM(2,0) TBATS KNN(5) COMB2.BG COMB2.SEL COMB4.SEL
SARIMAX(2,1,0)(1,0,0)24 - $<0.05$ $<0.05$ $<0.05$ $<0.05$ $<0.05$ $<0.05$ $<0.05$
FARX(3) 1.$0$ 1.0000 1.0000 1.0000 $<0.05$ 1.0000 1.0000
GAM(2,0) 1.0000 $<0.05$ 0.9191 0.0432 $<0.05$ $<0.05$ $<0.05$
TBATS 1.0000 $<0.05$ 0.0809 - 0.0001 $<0.05$ $<0.05$ $<0.05$
KNN(5) 1.0000 $<0.05$ 0.9568 0.$9$ $<0.05$ $<0.05$ $<0.05$
COMB2.BG 1.0000 1.0000 1.0000 1.0000 1.$0$ 1.0000 1.0000
COMB2.SEL 1.0000 $<0.05$ 1.0000 1.0000 1.0000 $<0.05$ 1.0000
COMB4.SEL 1.0000 $<0.05$ 1.0000 1.0000 1.0000 $<0.05$ $<0.05$
@@
SARIMA(1,1,0)(1,0,0)24 FARX(1) GAM(1,1) TBATS KNN(10) COMB2.BG COMB2.SEL COMB4.SEL COMB1.BG
SARIMA(1,1,0)(1,0,0)24 - $<0.05$ 1.0000 $<0.05$ 1.0000 $<0.05$ 1.0000 1.0000 $<0.05$
FARX(1) 1.$0$ 1.0000 0.9784 1.0000 1.0000 1.0000 1.0000 $<0.05$
GAM(1,1) $<0.05$ $<0.05$ $<0.05$ $<0.05$ $<0.05$ $<0.05$ $<0.05$ $<0.05$
TBATS 1.0000 $<0.05$ 1.$0$ 1.0000 0.8729 1.0000 1.0000 $<0.05$
KNN(10) $<0.05$ $<0.05$ 1.0000 $<0.05$ $<0.05$ 0.8906 0.9999 $<0.05$
COMB2.BG 1.0000 $<0.05$ 1.0000 0.1271 1.$0$ 1.0000 1.0000 $<0.05$
COMB2.SEL $<0.05$ $<0.05$ 1.0000 $<0.05$ 0.1094 $<0.05$ 1.0000 $<0.05$
COMB4.SEL $<0.05$ $<0.05$ 1.0000 $<0.05$ $<0.05$ $<0.05$ $<0.05$ $<0.05$
COMB1.BG 1.0000 0.9635 1.0000 1.0000 1.0000 1.0000 1.0000 1.$0$

### v2

@@ Appendix B Tables
SARIMAXa FARX(1) GAM(3,1) TBATS KNN(10) COMB3.SA COMB2.SEL
SARIMAXa - 1.00 <0.05<0.05 1.00 1.00 <0.05<0.05 <0.05<0.05
FARX(1) <0.05<0.05 - <0.05<0.05 1.00 1.00 <0.05<0.05 <0.05<0.05
GAM(3,1) 1.00 1.00 - 1.00 1.00 <0.05<0.05 <0.05<0.05
TBATS <0.05<0.05 <0.05<0.05 <0.05<0.05 - <0.05<0.05 <0.05<0.05 <0.05<0.05
KNN(10) <0.05<0.05 <0.05<0.05 <0.05<0.05 1.00 - <0.05<0.05 <0.05<0.05
COMB3.SA 1.00 1.00 1.00 1.00 1.00 - 1.00
COMB2.SEL 1.00 1.00 1.00 1.00 1.00 <0.05<0.05 -
  • Note. a the model order is (3,1,0)(1,1,1)24.
@@
SARIMAa FARX(3) GAM(3,1) TBATS KNN(10) COMB3.SA COMB2.SEL
SARIMAa - <0.05<0.05 <0.05<0.05 1.00 1.00 <0.05<0.05 <0.05<0.05
FARX(3) 1.00 - <0.05<0.05 1.00 1.00 <0.05<0.05 <0.05<0.05
GAM(3,1) 1.00 1.00 - 1.00 1.00 <0.05<0.05 <0.05<0.05
TBATS <0.05<0.05 <0.05<0.05 <0.05<0.05 - <0.05<0.05 <0.05<0.05 <0.05<0.05
KNN(10) <0.05<0.05 <0.05<0.05 <0.05<0.05 1.00 - <0.05<0.05 <0.05<0.05
COMB3.SA 1.00 1.00 1.00 1.00 1.00 - 1.00
COMB2.SEL 1.00 1.00 1.00 1.00 1.00 <0.05<0.05 -
  • Note. a the model order is (1,1,0)(1,1,1)24.
@@
SARIMAXa FAR(2) GAM(3,1) TBATS KNN(15) COMB3.SA COMB2.SEL
SARIMAXa - <0.05<0.05 <0.05<0.05 1.00 <0.05<0.05 <0.05<0.05 <0.05<0.05
FAR(2) 1.00 - <0.05<0.05 1.00 1.00 <0.05<0.05 <0.05<0.05
GAM(3,1) 1.00 1.00 - 1.00 1.00 <0.05<0.05 <0.05<0.05
TBATS <0.05<0.05 <0.05<0.05 <0.05<0.05 - <0.05<0.05 <0.05<0.05 <0.05<0.05
KNN(15) 1.00 <0.05<0.05 <0.05<0.05 1.00 - <0.05<0.05 <0.05<0.05
COMB3.SA 1.00 1.00 1.00 1.00 1.00 - 0.72
COMB2.SEL 1.00 1.00 1.00 1.00 1.00 0.28 -
  • Note. a the model order is (3,1,0)(1,1,1)24.
@@
SARIMAXa FARX(3) GAM(2,0) TBATS KNN(5) COMB3.SA COMB2.BG
SARIMAXa - <0.05<0.05 1.00 1.00 1.00 <0.05<0.05 <0.05<0.05
FARX(3) 1.00 - 1.00 1.00 1.00 <0.05<0.05 <0.05<0.05
GAM(2,0) <0.05<0.05 <0.05<0.05 - 0.92 0.04 <0.05<0.05 <0.05<0.05
TBATS <0.05<0.05 <0.05<0.05 0.08 - <0.05<0.05 <0.05<0.05 <0.05<0.05
KNN(5) <0.05<0.05 <0.05<0.05 0.96 1.00 - <0.05<0.05 <0.05<0.05
COMB3.SA 1.00 1.00 1.00 1.00 1.00 - 1.00
COMB2.BG 1.00 1.00 1.00 1.00 1.00 <0.05<0.05 -
  • Note. a the model order is (2,1,0)(1,1,1)24.
@@
SARIMAXa FARX(1) GAM(1,1) TBATS KNN(10) COMB1.BG COMB1.SA
SARIMAXa - <0.05<0.05 1.00 <0.05<0.05 1.00 <0.05<0.05 <0.05<0.05
FARX(1) 1.00 - 1.00 1.00 1.00 <0.05<0.05 <0.05<0.05
GAM(1,1) <0.05<0.05 <0.05<0.05 - <0.05<0.05 <0.05<0.05 <0.05<0.05 <0.05<0.05
TBATS 1.00 <0.05<0.05 1.00 - 1.00 <0.05<0.05 <0.05<0.05
KNN(10) <0.05<0.05 <0.05<0.05 1.00 <0.05<0.05 - <0.05<0.05 <0.05<0.05
COMB1.BG 1.00 1.00 1.00 1.00 1.00 - 1.00
COMB1.SA 1.00 1.00 1.00 1.00 1.00 <0.05<0.05 -
  • Note. a the model order is (1,1,0)(1,1,1)24.